OpenStudy (anonymous):

f(x)=5^x log4(x) The 4 is a subscript. Find the derivative. Would I need the power rule?

OpenStudy (anonymous):

Nope. Product rule, then use\[\frac{d}{dx}a^x=\ln a \cdot a^x\]and\[\frac{d}{dx}\log_b {x}=\frac{1}{x\ln b}.\]